Seth sold 16 tickets to a school play. Of these, a were adults' tickets at $3.75 each. The rest were children's tickets at $2.50 each. How much money did seth collect?
Answer by stanbon(75887)

Seth sold 16 tickets to a school play. Of these, a were adults' tickets at $3.75 each. The rest were children's tickets at $2.50 each. How much money did seth collect?
---------------
"a" adult tickets; value = 3.75a dollars
"16-a" child tickets; value = 2.5(16-a) dollars
EQUATION:
Seth collected 3.75a + (40-2.5a) = 1.5a + 40 dollars
----------------------
Cheers,
Stan H.